OpenStudy (anonymous):

Sheryl's mean average on eight exams is 83.750. Find the sum of her scores. A. 770 B. 670 C. 693 D. 686

OpenStudy (dumbcow):

avg = sum/n 83.75 = sum/8 sum = 8*83.75

OpenStudy (anonymous):

i get it. thanx

OpenStudy (anonymous):

I don't get it please help me understand what to do

OpenStudy (anonymous):

never mind I got it I think it;s B